Project

General

Profile

Actions

2017-07-27-LH2

Phone details

Notes

Actions

  • Action on CMcW to trace trigger signal to FC datalogger.
  • Action MT/JB to agree and implement testing of quench valves. In progress
  • Action PW to modify purge sequence limits in agreement with JB. In Progress
  • Action AK/PW to test permissions of required connections HMI to PLC.
    • Action CMcW/JB to search archiver for FC radiation shield temp & helium pressure excursions. Ongoing
  • Action PF/AK to propose IOC test program for next week. New expert panel has been tested - system details from PLC are not reported(can be removed) but information retrieved is correct. some additions have been agreed - implement tomorrow.
  • Action AK/JB to arrange further meeting on alarm handler development
  • Action JB/PF to agree wiki page for paperwork.
  • Action JB Atex rating of everything in H2 shed.
  • Action KL discuss with JT final paperwork required - note ISIS requires document giving clear instruction for ISIS MCR
  • Action JB to prepare rota on 'limited call-out' basis
  • Action CW to contact HR re 'on-call' times

Preliminary H2 tour

Actions recorded by AN.
  • Discussion and allocation

https://micewww.pp.rl.ac.uk/attachments/download/9080/Preliminary%20Safety%20tour%20of%20the%20MICE%20Liquid%20hydrogen%20installation.docx

Monitoring and archival of data - PF/AK.

  • IOC
  • Alarm handler

discuss possible additional archived variables beyond those for monitoring

Current technical

  • FC pressure rating.
  • Refurbishment of pipework from turret to wall.
  • Refurbishment of pipework from wall to roof and vent.

Neon Run

  • progress report PW/MT/HN

Shifters

  • Atex training Additonal staff for Atex on 20th 21st Sept.
    JB to discuss with DH H2 experts without Atex.

Timetable

Next meeting

  • Thurs 3rd August

AoB

Back to LH2:

Back

Files:

Updated by Whyte, Colin over 6 years ago · 3 revisions